In the year ended 31 December 20X9, there were 12 million ordinary shares in issue. In the year ended 31 December 20X9 the earnings available for ordinary shareholders amounted to INR 5 million. There are 1 million 10% INR 1 convertible loan notes in issue, convertible at the rate of 3 ordinary shares for every INR 4 of notes in the year ended 31/12/20X9 and the rate of company tax is 30%.

Required:

What is the fully diluted EPS for the year ended 31 December 20X9?
Answer #1

Calculation of Diluted EPS:-

1. Saving in Interest of loan :- INR (10,00,000 * 0.10) = 100,000

= INR 100,000 * (1 - 0.30 ) = 70,000

2. Increase in Number of Shares:-   INR 1,00,000 * 3/4 = 75,000

3. Diluted EPS :- ( 50,00,000+70,000 ) / (1,20,00,000 + 75,000)

= 50,70,000 / 1,20,75,000

= 0.42 ( Rounding up)